# Extreme events in locally coupled bursting neurons

### v1

@@ V Conclusion
The collective dynamics of the membrane potential of a network consisting of N unidirectionally coupled HR neurons, for a range of coupling strength, is found to spike occasionally to higher amplitudes that is $6$ times its standard deviation from the mean level. These rare spikes display dragon king-like statistics and a long tail in the Gaussian distribution substantiating that these are EE. Observing the two-parameter phase diagram between the coupling strength and input current, reveals the regions that exhibit EE. During the increase in the coupling strength, we also found that the collective observable transits from complete chaotic spiking to mixutre of chaotic spiking and bursting to complete bursting. During this transition, EE occur when the collective dynamics is either a chaotic spiking or a mixutre of chaotic spiking and bursting. During such a mixed dynamics, the EE occur as a result of intermittent cluster synchronization at the microscopic level. Due to the choice of ICs, odd neurons and even neurons tend to form separate clusters. These clusters are mostly out of phase with each other and only intermittently they are in phase synchronized state. During such intermittent phase synchronization, EE occur in the collective observable. Apart from this, increasing the number of different ICs increases the number of clusters in the system. For 100 different ICs, it can be observed that if 25-30 neurons get into a very short time zone, EE are produced. IEI plots and number of EE plot deciphers the rarity of the events for an increase in ICs.
Despite several mechanisms to EE are explored for lower dimensional networks, very little has been explored about the route to EE in higher dimensional coupled systems. Especially, no routes to EE in higher dimensional HR network has been determined so far. In this work, we have shown a new mechanism called the intermittent cluster synchronization as a new route of EE in unidirectionally and locally coupled HR neurons. Intermittent cluster synchronization biologically refers to the sudden synchronization occurring among the neurons in the brain, collectively leading to a detrimental disorder like, Epilepsy. Devising methods to stop this intermittent synchronization may help us to prevent a person from confronting epileptic seizures. Importantly, present result and the mechanism found out in this work serves as a milestone in the understanding of dynamical mechanisms behind epileptic seizures.
@@ Acknowledgements
Another kind of interaction commonly used to model the nearest neighbour interaction in network theory of coupled oscillators is the diffusive coupling. We analyze the emergence of EE in a unidirectionally coupled HR neurons interacting diffusively and the corresponding equation of motion is
@@
where $i=1,~{}2,~{}\ldots,~{}N$ is the number of neurons in the system with the boundary condition ($x_{N+1}=x_{1}$). The definition and values of the parameters are the same as in Sec. 2. These neurons are coupled in a ring form and look like asymmetrically coupled neurons. Similar to the chemical coupling case, in the diffusive coupling case also we observe the EE to occur in the system (LABEL:eqncoupled1) for various values of the coupling strength $k$. For the parameter space $k$, we plot the probability plot in Fig.  11.
Figure 11: The values of probability $P_{EE}$ of system (LABEL:eqncoupled1) by changing the coupling strength $k$ at $I=4$.
@@
In the inhibitory region, EE are found in the region from $k=-0.079$ to $k=-0.015$. Also in the excitatory regions, EE are found between $k=0.014$ and $k=0.072$. To visualize, we plot the time series of the collective observable for $k=-0.018$ in Fig. 12(a) where the trajectories cross the threshold $H_{s}=1.38$. The mechanism through which EE occur in the system (LABEL:eqncoupled1) is also intermittent cluster synchronization. To demonstrate this, in Fig. 12(b), we plot the spatio-temporal plot. Cluster synchronization can be seen by the bright green line in the spatio-temporal plot corresponding to the EE in the collective observable $X_{m}$.

### v2

@@ V Conclusion
The collective dynamics of the membrane potential of a network consisting of $N$ unidirectionally coupled HR neurons, for a range of coupling strength, is found to spike occasionally to higher amplitudes that is $6$ times its standard deviation from the mean level. These rare spikes display dragon king-like statistics and a long tail in the non-Gaussian distribution substantiating that these are EE. Observing the two-parameter phase diagram between the coupling strength and input current, reveals the regions that exhibit EE. During the increase in the coupling strength, we also found that the collective observable transits from complete chaotic spiking to mixture of chaotic spiking and bursting to complete bursting. During this transition, EE occur when the collective dynamics is either a chaotic spiking or a mixture of chaotic spiking and bursting. During such a mixed dynamics, the EE occur as a result of intermittent cluster synchronization at the microscopic level. Due to the choice of ICs, odd neurons and even neurons tend to form separate clusters. These clusters are mostly out of phase with each other and only intermittently they are in phase synchronized state. During such intermittent phase synchronization, EE occur in the collective observable. Apart from this, increasing the number of different ICs increases the number of clusters in the system. For 100 different ICs, it can be observed that if 25-30 neurons get into a very short time zone, EE are produced. IEI plots and number of EE plot deciphers the rarity of the events for an increase in ICs.
Despite several mechanisms to EE are explored for lower dimensional networks, very little has been explored about the route to EE in higher dimensional coupled systems. Especially, no routes to EE in higher dimensional HR network has been determined so far. In this work, we have shown a new mechanism called the intermittent cluster synchronization as a new route of EE in unidirectionally and locally coupled HR neurons. Intermittent cluster synchronization biologically refers to the sudden synchronization occurring among the neurons in the brain, collectively leading to a detrimental disorder like, epilepsy. Possible mitigation strategies like occasional random perturbations, pinning of oscillators, introducing other specific forms of coupling may serve as a new avenue in preventing this detrimental and unwanted event. Some possible Importantly, present result and the mechanism found out in this work serves as a milestone in the understanding of dynamical mechanisms behind epileptic seizures.
In this work, we observed EE in a symmetric network and intermittent cluster synchronization is observed irrespective on the symmetric in the initial conditions. In this connection, the symmetry breaking bifurcations in equivariant systems [52, 53, 54] could be an interesting problem in future to (i) identify the subgroup symmetries or equivariant bifurcations (e.g., Hopf or pitchfork) associated with the cluster states, and (ii) explore the emergence of anti-phase and in-phase synchronization as solutions constrained by the Cn symmetry group of the ring topology.
@@ Acknowledgements
Another kind of interaction commonly used to model the nearest neighbour interaction in network theory of coupled oscillators is the diffusive coupling. We analyze the emergence of EE in a unidirectional ring of diffusively coupled HR neurons. The corresponding equation of motion is
@@
where $i=1,~{}2,~{}\ldots,~{}N$ is the number of neurons in the system with the boundary condition ($x_{N+1}=x_{1}$). The definition and values of the parameters are the same as in Sec. 2. Similar to the chemical coupling case, in the diffusive coupling case also we observe the EE to occur in the system (LABEL:eqncoupled1) for various values of the coupling strength $k$. The probability of EE and bifurcation diagram with respect to coupling strength $k$ are shown in Figs.  14(a) and Fig. 14(b), respectively. Comparing both these figures, we can see that the whenever the system exhibits non-zero probability (Fig. 14(a)), correspondingly the size of the attractor is also high. Wherever the system is periodic, anti-phase synchronization occurs, resulting in the non-observance of extreme events. This happens throughout the parameter space $k$, except for higher values of the coupling strength (shown by blue dotted points in Fig.  14(b)) where burst synchronization occurs, and no extreme events appear in the system.
Figure 14: (a) The values of probability $P_{EE}$ and (b) bifurcation diagram of $X_{m}$ of system (LABEL:eqncoupled1) by varying coupling strength $k$ at $I=4$. Blue points in (b) represents chaotic regime with out extreme events. Here the periodic and chaotic regimes (red points) show non-EE and EE due to anti-phase and occasional in-phase synchronization between two clusters.
@@
In the repulsive coupling, EE are found in the region between $k=-0.079$ and $k=-0.015$. Also for the attractive coupling, EE are found between $k=0.014$ and $k=0.072$. To visualize, we plot the time series of the collective observable for $k=-0.018$ in Fig. 15(a) where the trajectories cross the threshold $H_{s}=1.38$. The mechanism through which EE occur in the system (LABEL:eqncoupled1) is also intermittent cluster synchronization. To demonstrate this, in Fig. 15(b), we plot the spatio-temporal plot. Cluster synchronization can be seen by the bright green line in the spatio-temporal plot corresponding to the EE in the collective observable $X_{m}$.
Figure 16: Bifurcation diagram of node-1 of system (LABEL:eqncoupled1) with respect to $I$. Here period-doubling route to chaos is observed by decreasing the value of $I$.
In Fig. 16, we show the dynamics of node 1 in the form of bifurcation diagram. We can observe that periodic nature interspersed by chaotic behaviour. This chaotic behaviour is observed by period doubling route.